OpenStudy (baldymcgee6):

If t = ln(5+/- 2 sqrt(6)) and u = ln(7+/- 4sqrt(3)), what are the possible values of coth(t) - tanh(u) ?

OpenStudy (baldymcgee6):

@AccessDenied, @mathslover

OpenStudy (baldymcgee6):

@Algebraic! do you have any insight to a solution?

OpenStudy (anonymous):

plug in the two values to coth and tanh? post results

OpenStudy (baldymcgee6):

here was the original question: If cosht=5 and coshu=7, what are the possible values of coth(t)−tanh(u) ? Do you think I am doing it right so far?

OpenStudy (baldymcgee6):

I solved cosht=5 and coshu=7 using quadratic formula

OpenStudy (anonymous):

looks fine

OpenStudy (baldymcgee6):

okay, when I sub in the t = ln(5+/- 2 sqrt(6)) and u = ln(7+/- 4sqrt(3)) values, I am not sure how to simplify it at all

OpenStudy (baldymcgee6):

Like obviously I could just throw that into the calculator, but I have to get the exact answer

OpenStudy (anonymous):

then use the definitions

OpenStudy (baldymcgee6):

right, so if i use the definitions, I dont know how to simplify it at all.

OpenStudy (anonymous):

\[e ^{\ln(x)} =?\]

OpenStudy (baldymcgee6):

ahhhhh

OpenStudy (baldymcgee6):

duh.. thanks

OpenStudy (anonymous):

:)

OpenStudy (baldymcgee6):

@Algebraic!, e^(-ln(x)) = ????

OpenStudy (anonymous):

1/x

OpenStudy (baldymcgee6):

hmm, okay. Thank you again!